Plz solve this Question!!!


The sum of the three numbers in A.P is 21 and the product of their extremes is 45. Find the numbers.

Grade:

1 Answers

raj kaushik
15 years ago

Let the numbers are be a - d, a, a + d

Then a - d + a + a + d = 21

3a = 21

a = 7

and (a - d)(a + d) = 45

a2 - d2 = 45

d2 = 4

d = +2

Hence, the numbers are 5, 7 and 9 when d = 2 and 9, 7 and 5 when d = -2. In both the cases numbers are the same.